# arXiv:2603.01511v1[cs.AI] 02 Mar 2026

### v1

@@ Introduction
To overcome the dual bottlenecks of scarce annotations and extreme class imbalance, current research has largely pursued three complementary paradigms: (i) harnessing sequence-only pre-training to transfer massive unlabeled evolutionary knowledge, as demonstrated by ESM-1b (Rives et al. 2021) and ProtTrans (Elnaggar et al. 2021); (ii) incorporating sparse but precise 3D cues via structure-aware encoders when experimental coordinates are available, such as MIF (Yang, Zanichelli, and Yeh 2023) and PST (Chen et al. 2024); and (iii) enhancing predictive capacity via cross-modal fusion that integrates auxiliary modalities like protein textual descriptions through sequence-text embedding alignment, including MMSite (Ouyang et al. 2024), UniSite (Fan et al. 2025), and ProtST (Xu et al. 2023). Despite substantial progress, two persistent challenges persist in achieving reliable active-site prediction.
@@
To address these challenges, we introduce Multimodal Mixture-of-Experts with Retrieval Augmentation (MERA), a novel framework for protein active site identification. Our approach incorporates two key innovations. Multi-expert RAG (MeRAG) addresses the vulnerability of single-instance prediction by extracting meaningful information from sparse retrieval sequences through three orthogonal retrieval experts (chain, sequence, and active-site). Each expert captures distinct biological information, and their outputs are dynamically aggregated via a residue-level Mixture-of-Experts (MoE) gating mechanism. This design enables adaptive and robust feature learning even under sparse data conditions. (ii) Reliability-aware Multimodal Fusion (RMF) addresses inadequate modality reliability estimation by explicitly quantifying the trustworthiness of each modality. Inspired by Dempster–Shafer evidence theory (Amini et al. 2020; Huang et al. 2025; Deregnaucourt et al. 2025), we model the prediction of each modality as a belief mass function and apply learnable discounting coefficients to reflect modality reliability. This principled approach enables the model to appropriately attenuate less reliable modalities during the fusion process, offering more robust multi-modal integration when information quality varies across modalities (Han, Chen, and Ban 2024). Guided by these reliability scores, we perform residue-level fusion to achieve more trustworthy predictions.
@@
Protein active sites are spatially clustered residues within proteins that are directly responsible for molecular recognition (Nussinov 2025), substrate binding (Ji et al. 2024) and catalyzing enzymatic reactions (Reisenbauer, Sicinski, and Arnold 2024). Accurate identification of these functionally critical regions is fundamental to mechanistic biology and therapeutic design, as they directly determine protein structural and functional roles. Existing methods for protein active site identification can be broadly categorized into single-modality and multi-modal approaches.
@@
form the foundation of most current work and encompass both sequence-based and structure-based methods. Sequence-based models such as ESM-1b (Rives et al. 2021), ESM-1v (Meier et al. 2021), ESM-2 (Lin et al. 2022), and ProtTrans (Elnaggar et al. 2021) learn residue representations directly from amino acid sequences through masked language modeling, demonstrating strong generalization across diverse functional prediction tasks. Structure-based models utilize explicit 3D geometric information, typically employing graph neural networks or geometric encoders such as MIF (Yang, Zanichelli, and Yeh 2023) and PST (Chen et al. 2024), to improve structure-dependent site prediction. However, both approaches are inherently limited by relying on information from a single modality.
@@
where $m^{s}_{i}$ the relative evidence strength of modality $s$ for residue $i$.
To assess the modality trustworthiness beyond simple prediction confidence, we compute a credibility (discounting) coefficient following (Han, Chen, and Ban 2024):
@@
where lower $u^{s}_{i}$ values indicate higher reliability for modality $s$ at residue $i$.
@@
where $e^{s}_{i}$ is the evidence-based fusion weight for modality $s$ at residue $i$. The final prediction is computed as a reliability-weighted combination of the logits from all modalities:
@@
  • Active site identification (ProTAD-Gen): we compare against sequence-only models, including ESM-1b (Rives et al. 2021), ESM-1v (Meier et al. 2021), ESM-2 (Lin et al. 2022), ProtElectra (Elnaggar et al. 2021), PETA (Tan et al. 2024), TAPE (Rao et al. 2019), and S-PLM (Wang et al. 2025); structure-aware models PST (Chen et al. 2024) and MIF (Yang, Zanichelli, and Yeh 2023); and multi-modal models MMSite (Ouyang et al. 2024), ProtST (Xu et al. 2023), and UniSite (Fan et al. 2025).
  • Peptide-binding site prediction (TS125): We evaluate against sequence-based deep models IIDL-PepPI (Chen et al. 2025), PepCA (Huang et al. 2024), PepBCL (Wang et al. 2022), PepCNN (Chandra et al. 2023), and PepNN (Abdin et al. 2022); protein language models ESM-2 (Lin et al. 2022) and TAPE (Rao et al. 2019); structure-aware methods: PepBind (Zhao, Peng, and Yang 2018) and PepSite (Petsalaki et al. 2009).
@@
Protein sequences are encoded using ESM-1b (1280-dimensional embeddings), while textual descriptions are processed by BioMedBERT (768-dimensional embeddings). For inference, the original MMSite framework requires UniProt IDs to generate protein descriptions via Prot2Text. To enable broader applicability, particularly for novel proteins lacking UniProt annotations, we replace this with sequence-to-text models like ESM2Text, which directly generates functional annotations solely from raw sequence data. For retrieval, we select the top-$3$ nearest neighbors within each cluster, achieving an effective balance between computational efficiency and retrieval performance, as validated by our ablation studies. All experiments are conducted on two 40 GB NVIDIA A100 GPUs. Models are trained for 100 epochs, requiring approximately three hours for each task. For the TS125 dataset, we extend our framework with an additional peptide expert. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$. Our model contains approximately 300M parameters.

### v2

# Multimodal Mixture-of-Experts with Retrieval Augmentation for Protein Active Site Identification

@@ Introduction
To overcome the dual bottlenecks of scarce annotations and extreme class imbalance, current research has largely pursued three complementary paradigms: (i) harnessing sequence-only pre-training to transfer massive unlabeled evolutionary knowledge, as demonstrated by ESM-1b (Rives et al. 2021) and ProtTrans (Elnaggar et al. 2021); (ii) incorporating sparse but precise 3D cues via structure-aware encoders when experimental coordinates are available, such as MIF (Yang et al. 2023) and PST (Chen et al. 2024); and (iii) enhancing predictive capacity via cross-modal fusion that integrates auxiliary modalities like protein textual descriptions through sequence-text embedding alignment, including MMSite (Ouyang et al. 2024), UniSite (Fan et al. 2025), and ProtST (Xu et al. 2023). Despite substantial progress, two persistent challenges persist in achieving reliable active-site prediction.
@@
To address these challenges, we introduce Multimodal Mixture-of-Experts with Retrieval Augmentation (MERA), a novel framework for protein active site identification. Our approach incorporates two key innovations. Multi-expert RAG (MeRAG) addresses the vulnerability of single-instance prediction by extracting meaningful information from sparse retrieval sequences through three orthogonal retrieval experts (chain, sequence, and active-site). Each expert captures distinct biological information, and their outputs are dynamically aggregated via a residue-level Mixture-of-Experts (MoE) gating mechanism. This design enables adaptive and robust feature learning even under sparse data conditions. (ii) Reliability-aware Multimodal Fusion (RMF) addresses inadequate modality reliability estimation by explicitly quantifying the trustworthiness of each modality. Inspired by Dempster–Shafer evidence theory (Amini et al. 2020; Huang et al. 2025; Deregnaucourt et al. 2025), we model the prediction of each modality as a belief mass function and apply learnable discounting coefficients to reflect modality reliability. This principled approach enables the model to appropriately attenuate less reliable modalities during the fusion process, offering more robust multi-modal integration when information quality varies across modalities (Han et al. 2024). Guided by these reliability scores, we perform residue-level fusion to achieve more trustworthy predictions.
@@
Protein active sites are spatially clustered residues within proteins that are directly responsible for molecular recognition (Nussinov 2025), substrate binding (Ji et al. 2024) and catalyzing enzymatic reactions (Reisenbauer et al. 2024). Accurate identification of these functionally critical regions is fundamental to mechanistic biology and therapeutic design, as they directly determine protein structural and functional roles. Existing methods for protein active site identification can be broadly categorized into single-modality and multi-modal approaches.
@@
form the foundation of most current work and encompass both sequence-based and structure-based methods. Sequence-based models such as ESM-1b (Rives et al. 2021), ESM-1v (Meier et al. 2021), ESM-2 (Lin et al. 2022), and ProtTrans (Elnaggar et al. 2021) learn residue representations directly from amino acid sequences through masked language modeling, demonstrating strong generalization across diverse functional prediction tasks. Structure-based models utilize explicit 3D geometric information, typically employing graph neural networks or geometric encoders such as MIF (Yang et al. 2023) and PST (Chen et al. 2024), to improve structure-dependent site prediction. However, both approaches are inherently limited by relying on information from a single modality.
@@
where $m^{s}_{i}$ the relative evidence strength of modality $s$ for residue ii.
To assess the modality trustworthiness beyond simple prediction confidence, we compute a credibility (discounting) coefficient following (Han et al. 2024):
@@
where lower $u^{s}_{i}$ values indicate higher reliability for modality $s$ at residue ii.
@@
where $e^{s}_{i}$ is the evidence-based fusion weight for modality $s$ at residue ii. The final prediction is computed as a reliability-weighted combination of the logits from all modalities:
@@
  • Active site identification (ProTAD-Gen): we compare against sequence-only models, including ESM-1b (Rives et al. 2021), ESM-1v (Meier et al. 2021), ESM-2 (Lin et al. 2022), ProtElectra (Elnaggar et al. 2021), PETA (Tan et al. 2024), TAPE (Rao et al. 2019), and S-PLM (Wang et al. 2025); structure-aware models PST (Chen et al. 2024) and MIF (Yang et al. 2023); and multi-modal models MMSite (Ouyang et al. 2024), ProtST (Xu et al. 2023), and UniSite (Fan et al. 2025).
  • Peptide-binding site prediction (TS125): We evaluate against sequence-based deep models IIDL-PepPI (Chen et al. 2025), PepCA (Huang et al. 2024), PepBCL (Wang et al. 2022), PepCNN (Chandra et al. 2023), and PepNN (Abdin et al. 2022); protein language models ESM-2 (Lin et al. 2022) and TAPE (Rao et al. 2019); structure-aware methods: PepBind (Zhao et al. 2018) and PepSite (Petsalaki et al. 2009).
@@
Protein sequences are encoded using ESM-1b (1280-dimensional embeddings), while textual descriptions are processed by BioMedBERT (768-dimensional embeddings). For inference, the original MMSite framework requires UniProt IDs to generate protein descriptions via Prot2Text. To enable broader applicability, particularly for novel proteins lacking UniProt annotations, we replace this with sequence-to-text models like ESM2Text, which directly generates functional annotations solely from raw sequence data. For retrieval, we select the top-33 nearest neighbors within each cluster, achieving an effective balance between computational efficiency and retrieval performance, as validated by our ablation studies. All experiments are conducted on two 40 GB NVIDIA A100 GPUs. Models are trained for 100 epochs, requiring approximately three hours for each task. For the TS125 dataset, we extend our framework with an additional peptide expert. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$. Our model contains approximately 300M parameters.